# Geo-Sign: Hyperbolic Contrastive Regularisation for Geometrically Aware Sign Language Translation

### v1

@@ Appendix G Code Listings
1class GCNUnit(nn.Module):
@@
5 self.A = nn.Parameter(A.clone()) if adaptive else A
6 self.conv = nn.Conv2d(Cin, Cout*self.K, (1,1))
7 self.bn = nn.BatchNorm2d(Cout)
8 self.act = nn.ReLU(inplace=True)
@@
16class STGCNBlock(nn.Module):
@@
21 self.tcn = nn.Sequential(
22 nn.Conv2d(Cout, Cout, (t_kernel,1), (stride,1), (pad,0)),
23 nn.BatchNorm2d(Cout))
24 self.res = (nn.Identity() if Cin==Cout and stride==1
25 else nn.Conv2d(Cin, Cout, 1, (stride,1)))
26 self.act = nn.ReLU(inplace=True)
@@
1class HyperbolicProjection(nn.Module):
@@
5 self.proj = nn.Linear(d_in, d_out)
6 self.log_scale = nn.Parameter(torch.zeros(())) # ln(scale)
@@
1class HyperbolicContrastiveLoss(nn.Module):
@@
5 self.log_tau = nn.Parameter(torch.logit(torch.tensor(tau0/2)))
6 self.m = nn.Parameter(torch.tensor(m0))

### v2

@@ Appendix G Code Listings
1class GCNUnit($n$.Module):
@@
5 self.A = $n$.Parameter(A.clone()) if adaptive else A
6 self.conv = $n$.Conv2d(Cin, Cout*self.K, (1,1))
7 self.bn = $n$.BatchNorm2d(Cout)
8 self.act = $n$.ReLU(inplace=True)
@@
16class STGCNBlock($n$.Module):
@@
21 self.tcn = $n$.Sequential(
22 $n$.Conv2d(Cout, Cout, (t_kernel,1), (stride,1), (pad,0)),
23 $n$.BatchNorm2d(Cout))
24 self.res = ($n$.Identity() if Cin==Cout and stride==1
25 else $n$.Conv2d(Cin, Cout, 1, (stride,1)))
26 self.act = $n$.ReLU(inplace=True)
@@
1class HyperbolicProjection($n$.Module):
@@
5 self.proj = $n$.Linear(d_in, d_out)
6 self.log_scale = $n$.Parameter(torch.zeros(())) # ln(scale)
@@
1class HyperbolicContrastiveLoss($n$.Module):
@@
5 self.log_tau = $n$.Parameter(torch.logit(torch.tensor(tau0/2)))
6 self.m = $n$.Parameter(torch.tensor(m0))